# Clear Minds Think Alike: What Makes LLM Fine-tuning Robust? A Study of Token Perplexity

### v1

@@ 4.1 In Domain and Out Domain Datasets
We focus on three domains—Programming, Mathematics, and Knowledge-based tasks—to study performance shifts after supervised fine-tuning (SFT) with LoRA on each domain. For the Programming and Mathematics domains, we select MBPP and MATH for training, while others and Knowledge-based tasks for out-of-domain evaluation to assess generalization capabilities.
@@ 4.2 Models
We evaluate our method across different model scales and architectures using three publicly available instruction-tuned language models; here they are referred to as Base Models:
@@
  • Gemma2-2B Instruct: A 2.6 billion parameter model designed for resource-efficient deployment
  • Mistral-7B v0.3: A 7 billion parameter model featuring grouped-query attention and sliding window attention mechanisms.
@@
  • Llama3-8B Instruct: An 8 billion parameter model which has the best overall performance in all 3 models.
@@
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama3-8B and Mistral-7B represent mid-scale models with different architectural innovations, while Gemma2-2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies.
5 STM Experiement Results
@@
Table 2 shows that by setting a threshold to filter out high perplexity tokens as 2.5, which approximately filtering out around 20% to 24% of total tokens, with original instruct model on the ground truth dataset as training data, the testing performance is comparable to the Self-Output method. This strategy not only maintains in-domain performance but also preserves out-of-domain generalization. This result agrees with our hypothesis that the presence of high perplexity tokens is one of the causes of performance degradation of ground truth model, and low perplexity tokens sufficiently mitigate such negative impact for strong generalization without the need for self-generated training data.
Table 3: STM performance comparison between Gemma2-2B Instruct and Llama3-8B Instruct. The percentage indicates the number of filtered tokens from the training data.
Model MATH MBPP GSM8K ARC BIRD
Gemma2-2B Instruct
Ground Truth 21.7 45.8 19.0 29.5 5.1
$STM_{\tau=1000}$ (2.5%) 23.3 48.9 24.9 25.4 12.8
$STM_{\tau=25}$ (9.3%) 27.0 48.2 49.7 72.9 13.1
$STM_{\tau=2.5}$ (23.8%) 30.3 49.5 55.4 75.3 13.3
Llama3-8B Instruct
Ground Truth 25.8 58.7 48.6 67.9 19.4
$STM_{\tau=1000}$ (1.7%) 27.7 59.2 69.3 71.2 19.4
$STM_{\tau=10}$ (11.0%) 33.0 61.1 77.6 75.6 19.5
$STM_{\tau=2.5}$ (22.2%) 33.4 61.6 75.1 75.1 20.3
@@ 5.1 Optimal threshold selection for STM
A key hyperparameter of STM is the perplexity threshold to filter tokens from the supervised fine-tuned (SFT) model. We explore the existence of an optimal threshold that maximizes performance. Figure 3 reveals that filtering approximately 24% of tokens yields the best results. This optimal threshold demonstrates consistent benefits across both in-domain and out-of-domain datasets, including GSM8K, ARC, and MBPP, showing its robustness and generalizability similar to Self-Output data.
Figure 3: MATH SFT using STM method on Llama3-8B Instruct on different levels of token filtering levels. The best in domain performance nearly matches peak performance on out of domain tasks : GSM8k, ARC, MBPP as well.
5.2 Alternative Token Selection Strategies
In addition to that STM uses initial model perplexity for token selection, we investigate an alternative strategy.
Cross-scale Filtering: We explore perplexity assessment using larger models within the same model family (e.g., Gemma2-9B Instruct) to guide token selection for smaller models (e.g., Gemma2-2B Instruct), as denoted as ${STM}_{9B\tau=2.5}$. This approach investigates whether token selection benefits from a more robust understanding of larger models, potentially offering a form of knowledge distillation through perplexity-based filtering.
Table 12 shows the results for the best-performance models of the original STM and the alternative of STM. Compared to base model performance, both STM and $STM_{9B}$ outperform the base model results in term of in-domain and out-of-domain tasks, showing STM’s effectiveness in larger scaled models still benefits model training. However, even though selecting similar filter rate, scaling up perplexity model does not outperform the original STM. It is likely that a more capable model could score with lower perplexity on tokens that smaller-sized models perceive as high perplexity tokens. Therefore, to optimize the performance of STM, our token selection strategy is preferable to apply the same model as the training model, and set a threshold to match an ideal filtering rate(around 20% to 25%). We also investigated other alternatives, please see Appendix D.
Table 4: List of alternative token selection methodology supervised fine-tuned on MATH in Gemma2-2B Instruct. Here we choose Gemma2 series because Gemma2 has feasible model sizes to implement the cross-scale setting for comparison.
Gemma2-2B Filter MATH MBPP GSM8K ARC BIRD
Base model 0% 28.1 51.1 58.5 74.7 13.9
Ground Truth 0% 21.7 45.8 19.0 29.5 5.1
$STM$ 23.8% 30.3 49.5 55.4 75.3 13.3
${\tau=2.5}$
${STM}_{9B}$ 21.4% 27.2 47.6 42.0 72.4 14.1
${\tau=2.0}$
6 Additional Analysis Results
Other than STM experiment results, we conducted additional analysis on the learning and robustness of our mentioned methods.
6.1 Does Self-Output Response Correctness Matters?
In our experiment, the lowest perplexity scored data are collected exclusively from correct Self-Output subsets. In MBPP training task, a correct Self-Output data refers to a generated response that passes all the test cases. In MATH, correctness refers to a correct final answer parsed from a response. Hence, it raises an important research question regarding the necessity of correctness criteria when simultaneously pursuing in-domain performance improvements and minimizing out-of-domain degradation. In Table 5, we collect different correctness rates of Self-Output response as MBPP training datasets respectively with Llama3-8B Instruct model. Correctness rate refers to the ratio of the samples whose labels are verified as "correct". Our results show that the correctness of MBPP training data strongly affects the in-domain performance. It is intuitive that training with incorrect answers leads to lower in-domain performance. However, compared to model trained on ground truth data (the correctness rate is 100%), which fails to sustain out-of-domain task performance, models trained with higher correctness rate of Self-Output data achieves higher in-domain improvements and still sustains the out-of-domain task performance robustness at the same time. Therefore, high correctness improves in-domain performance, while Self-Output data effects out-of-domain performance more than correctness. We also have the correctness test with smaller models like Gemma2-2B Instruct, please refer to Appendix E.
Table 5: The performance of models trained with different correctness rate of MBPP Self-Output dataset and trainable tokens number. Values of a cell block in gray represents the in-domain testing performance of the model trained with in-domain task data.
rejection rate Trained tokens MBPP MATH ARC GSM8k BIRD
GT 75,464 .579 .233 .591 .185 .176
Self-output
100% 65,602 .606 .324 .793 .773 .199
75% 69,605 .593 .346 .794 .778 .199
0% 59,738 .561 .315 .802 .770 .188
base model - .593 .314 .782 .726 .206
6.2 Convergence Performance Across Different Methods
Training low perplexity data also leads to a low training loss intuitively, and model could converge within fewer training epochs. Figure 4 showcases the difference of training procedure on the ground truth data and the Self-Output data on MBPP task. The results show that training with Self-Output data converges as a lower validation loss curves within half of the training steps (3 epochs v.s. 7 epochs) compared to training with ground truth data. Moreover, within small steps of training (3 epochs), there is a higher testing performance of Self-Output data training (improved $+1.41\%$ compared to original Llama) than the performance of ground truth data training (degraded $-6.62\%$). STM method also shows the similar trends. This analysis suggests that it is more efficient to train with low perplexity data to achieve higher in-domain performance than to train with high perplexity data.
@@ 6.2 Convergence Performance Across Different Methods
Figure 4: MBPP in-domain testing accuracy, validation and training loss of ground truth, Self-Output, and STM data with 2.5 as perplexity filtering threshold.
6.3 Can Regularization Methods Train Robust Models?
We examine alternative regularization techniques for mitigating out-of-domain performance degradation during in-domain fine-tuning. Our comparative analysis encompasses two widely-adopted regularization approaches: Dropout [25] and Weight Decay [13]. These methods have demonstrated effectiveness in preventing model from overfiting. Table 6 shows that tuning weight decay and dropout rate is not a robust approach to achieving in-domain performance improvement and avoiding out-of-domain performance degradation at the same time, while Self-Output data training simply improves both. For more detail about the experiment settings and results, we leave them in appendix F.
Table 6: Comparison between Self-output model, Base model and Ground Truth models applying different combinations of regularization parameters. *Note that the regularization parameters of the highest MBPP and MATH performances are not the same.
Model Self Base Regular-
finetuned on MBPP Out Model ization
MBPP .616 .598 .558*
MATH .334 .314 .329*
6.4 LoRA Weight Analysis
@@
We also investigate the relationship between token level perplexity and the characteristics of weight updates in Low-Rank Adaptation (LoRA). Our analysis focuses on understanding how different types of training data affect the magnitude and dimensionality of parameter adjustments in transformer-based language models. Given that LoRA effectively learns offset weights $\Delta W$ to the original model parameters $W$, we calculate the L2 norm of $\Delta W$ to quantify the degree of deviation introduced by each fine-tuning dataset. As shown in Table 7, fine-tuning with Self-Output data consistently results in smaller L2 norm of weight updates compared to Rephrase and Ground Truth data, while STM produces the smallest updates (0.45-0.56) across all models. These minimal perturbations suggest that both approaches, particularly STM, help preserve the model’s original capabilities.
@@
Table 7: L2 norm of $\Delta W$ larger value signifies a larger update from the original weight $W$
Model Self Re- Ground STM
finetuned on MBPP Out phrase Truth
Llama3-8B Instruct 6.53 7.31 17.75 0.55
Gemma2-2B Instruct 4.03 5.78 5.69 0.45
@@ 7 Related Work
Instruction following [21, 3] is widely used to align the responses of LLMs with target task values. Practically, instruction following requires instruction tasks dataset to fine-tune LLMs, where each dataset consists of instructions and desired responses. Hence, the performance of instruction-tuning heavily relies on the quality of instruction data such as context richness [33, 29, 37]. In addition, Ghosh et al. [7] investigated the limitation of instruction-tuning of LLMs about the performance degradation (as known as catastrophic forgetting) from pattern-copying behaviors and hallucinations with Lora/full fine-tuning. These works give us a good start on how performance degrades in terms of models’ response behaviors and benchmarks when fine-tuning with instruction following dataset.
@@
As [29, 3] shows, since LLM breakdown easily after training with different tasks. There are several remedies to the performance improvement focusing on training data augmentation [30, 34]. For instance, [23] use much larger LLM such as GPT-4 and Claude to generate responses of questions as training labels, which improves both the target task and others non-target tasks performances. However, such distillation method neglects the correctness of generated labels for training data so that incorrect responses could be generated as the amount of generated data increases, and using equal or smaller size models for distillation could be challenging. [35] prompts LLMs to simply rephrase response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing ground truth answer limit the output distribution and result in lower performance in our study. Furthermore, [9] exploits a base LLM as a judge to pick out answerable questions and unanswerable questions to compose a new training dataset to improve in-domain performance only on QA/conversation dataset. Although using Mistral-7b v0.2 to generate acceptable responses for answerable question improves in-domain and out-of-domain tasks, it is rarely discussed that if the proposed method is applicable for different model sizes and series. Besides, using LLM as a judge still brings noise to the correctness of training, which shows to be important for in-domain training in our study. Lastly, [15, 17] propose a Selective Language Modeling (SLM) to score favorable tokens from training data, and training on selected tokens could bring higher performance. However, the requirements for pre-training a reference model or prepare a high-quality dataset for scoring model could be exhausting depending on task difficulty. Our proposed STM does not require such training before selection of tokens, which brings both efficiency and performance improvement at the same time.
@@ 8 Conclusion
In this work, we present a possible explanation for the effectiveness of fine-tuning on LLM outputs. Our analysis and empirical result reveal that the improved robustness stems from reduced high perplexity tokens due to Self-Output sequences. We proposed an alternative method, STM, which filters out high perplexity tokens from ground truth data. We conducted extensive experiments across Gemma2-2B, Mistral-7B and Llama3-8B, and the results showed that the STM method is competitive with Self-Output data in both in-domain and out-of-domain tasks. Our findings suggest that token-level perplexity plays an important role in the fine-tuning process, offering future robust fine-tuning a simple but effective baseline.

### v2

@@ 4.1 Target and non-target datasets
We focus on three domains, programming, mathematics, and knowledge-based tasks, to study performance degradation after supervised fine-tuning with LoRA on each domain. For the programming and mathematics domains, we select MBPP and MATH for target task training, while others and knowledge-based tasks for non-target task evaluation to assess generalization capabilities.
@@ 4.2 Models
  • Gemma 2 IT 2B: A 2.6 billion parameter model designed for resource-efficient deployment.
@@
  • Llama 3 8B Instruct: An 8 billion parameter model which has the best overall performance in all models.
@@
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama 3 8B Insturct represent mid-scale models with different architectural innovations, while Gemma 2 IT 2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies. To further test the generalization of STM effectiveness on a brand new model we also choose three additional new models to test on target task MBPP and MATH, please refer to Appendix B.
5 STM Results
@@
Table 2 shows that with a threshold to filter out high perplexity tokens such as 2.5, which approximately filters out around 20% to 24% of total tokens, its BWT is close to $0\%$, and STM also achieves comparable TI to the Self-Output method. It indicates that STM not only improves target task performance but also reduces non-target task degradation. This result agrees with our hypothesis that the presence of high perplexity tokens is one of the causes of performance degradation of the ground truth model, and low perplexity tokens sufficiently mitigate such negative impact for strong generalization without the need for self-generated data. We also applied STM to other models, and the results can be found in Appendix B. Our conclusions are still the same across different models.
5.1 Does filtering high perplexity tokens always perform better?
From the intuition of the high perplexity filtering mechanism of STM, we have done an ablation study on the masking criteria of perplexity by filtering out high, random, and low ppl tokens. More specifically, high ppl refers to filtering the top 25% ppl of all tokens, random filtering refers to filtering the randomly selected 25% of all tokens, and low ppl refers to filtering the lowest 25% ppl of all tokens from the dataset. As the upper part of Table 4 and Table 4 show, masking high ppl tokens in STM simultaneously reduces performance degradation and improves target task performance. Likewise, we have the same conclusions for Llama 3 8B Instruct on such ablation test, please refer to Appendix B.
Table 3: Ablation and scaling on threshold of STM with Gemma 2 IT 2B on MBPP task. Percentages indicate the ratio of tokens filtered from training data. STM performs generally the best with 20% to 24% of high ppl tokens filtered.
BWT(%) TI(%)
STMτ=2.5,high   0.4   0.0
STMτ=2.5,random -8.6 -15.6
STMτ=2.5,low -7.9 -18.7
Baseline Fine-Tuning -38.2 -25.2
STMτ=1000 (6.26%) -2.9 -11.4
STMτ=25 (12.34%) -2.5 -8.8
STMτ=10 (15.1%) -0.7 -10.4
STMτ=2.5 (23.8%)   0.4   0.0
STMτ=1.5 (26.1%) -0.3 -0.5
STM9Bτ=2.5 (23.8%) -3.8 -7.3
Table 4: Ablation and scaling on threshold of STM with Llama3 8B Instruct on MBPP and MATH task respectively. Percentages indicate the ratio of tokens filtered from training data. STM performs generally the best with 20% to 24% of high ppl tokens filtered.
BWT(%) TI(%)
STMτ=2.5,high   0.16   3.12
STMτ=2.5,random -6.64 -9.84
STMτ=2.5,low -9.15 -14.73
$STM_{\tau=1000}$ (1.7%) -19.6 -11.8
$STM_{\tau=10}$ (11.0%) -13.1   5.1
$STM_{\tau=2.5}$ (22.2%)   0.16   3.12
Table 5: STM applied on full weight fine-tuning (FWFT), DoRA and LoRA with MBPP on Gemma 2 IT 2B. STM can enhance all three kinds of fine-tuning techniques with much better preservation of non-target task capability and target task improvement.
[…]
FWFT -31.87 -27.98
FWFT + STMτ=2.5 -0.13 -8.81
LoRA -21.76 -38.19
LoRA + STMτ=2.5   0.42   0.0
DoRA -8.54 -15.2
DoRA + STMτ=2.5 -0.01 -0.04
5.2 Is larger-sized LLM a better token filter?
In addition to that, STM uses initial model perplexity for token selection, we investigate an alternative token selection model for STM.
Cross-scale Filtering. We explore perplexity assessment using larger models within the same model family (e.g., Gemma 2 IT 9B) to guide token selection for smaller models (e.g., Gemma 2 IT 2B), denoted as STM9Bτ=2.5. This approach investigates whether token selection benefits from a more robust understanding of larger models, potentially offering a form of knowledge distillation through perplexity-based filtering.
Table 4 shows that both STM and STM9B perform better than baseline fine-tuning. However, scaling the perplexity model does not surpass STM with masking created by its own (2B), likely because larger models assign lower perplexity to tokens smaller models find challenging. Therefore, the optimal strategy should always be to use the same model for token selection. We also investigated other alternatives such as a learnable token selection, please see Appendix E.
@@ 5.3 Optimal threshold selection for STM
A key hyperparameter of STM is the perplexity threshold to filter tokens from the supervised fine-tuned (SFT) model. We explore the existence of an optimal threshold that maximizes performance. The bottom part of Table 4 reveals that filtering approximately 24% of tokens yields the best results. This optimal threshold demonstrates consistent benefits across both target and non-target datasets, including GSM8K, ARC, MATH, and BIRD, showing its robustness and generalization similar to Self-Output data. For further validation of the generalization of such settings, please refer to Table 9 in Appendix B that such settings apply to different model families as well.
5.4 STM generalization on different fine-tuning strategies
As STM is a masking technique applied to target data during the fine-tuning stage, it should be orthogonal to the fine-tuning techniques and improve the performance degradation. Thus, we experiment with how STM improves practical fine-tuning strategies like Full-Weight Fine-tuning (FWFT), Low-rank adaptation (LoRA), and Weight-decomposed LoRA (DoRA). As Table 5 shows, all of the fine-tuning techniques are further improved by STM, which indicates STM’s generalization on different fine-tuning strategies. We have the same conclusions for Llama 3 8B Instruct with positive results on LoRA and DoRA in Appendix B.
5.5 A continual learning setting of STM
We have experimented the effectiveness of STM by training on a single target task to reduce non-target tasks degradation. Yet, in a setting of continual learning scenario as [14] presents, degradation of tasks also occurs after several target tasks are trained sequentially. We simulate such setting to fine-tune Llama 3 8B Instruct with MATH initially and followed by MBPP tasks. Table 6 shows that (1) not only the target task is improved after fine-tuning, but also its degradation is very small after training another target task; (2) the non-target tasks like GSM8K, ARC and BIRD keeps similar performance even after two rounds of fine-tuning. It indicates STM can be effective in such continual learning setting as well.
Table 6: STM applied on continual learning settings
steps Training Task MATH-test MBPP-test GSM8K ARC BIRD
0 Initial model 0.3140 0.5979 0.7255 0.7816 0.2053
1 MATH-train 0.334 0.6164 0.7687 0.7577 0.2027
2 MBPP-train 0.332 0.625 0.7635 0.7688 0.2073
6 Analysis and Discussion
In our experiments, we found that STM’s effectiveness is applicable across different datasets, models, and fine-tuning strategies. In this section, we conduct several analyses on the model training process and model changes with STM to investigate the reason that reduces performance degradation.
Low training loss from low perplexity training
Training low perplexity data leads to a low training loss intuitively, and the model can converge within fewer training epochs. Figure 4 showcases the training procedure on MBPP training data of different perplexities. The results show that training with STM or synthesized data is fit with much lower training loss, resulting in fewer updates of model weight for changing capabilities of non-target tasks.
@@ 6 Analysis and Discussion
Figure 4: MBPP target task testing accuracy, validation and training loss of baseline finetuning with ground truth data, finetuning with Self-Output data, and STM strategy of perplexity filtering threshold=2.5 for Llama 3 8B Instruct. STM and self-output training yield better performances with much lower training and validation loss because of low-perplexity training.
Subtle LoRA weight changed from low perplexity training
@@
We investigate the relationship between token-level perplexity and characteristics of weight updates in LoRA. Our analysis focuses on understanding how different types of training data affect the magnitude and dimensionality of parameter adjustments in transformer-based language models. Given that LoRA effectively learns offset weights $\Delta W$ to the original model parameters $W$, we calculate the L2 norm of $\Delta W$ to quantify the degree of deviation introduced by each fine-tuning dataset. As shown in Table 7, fine-tuning with Self-Output data consistently results in smaller L2 norm of weight updates compared to Rephrase and Ground Truth data, while STM produces the smallest updates (0.45-0.56) across all models. Such minimal perturbations suggest that both approaches, particularly STM, help preserve the model’s original capabilities. Please see Appendix H for details on the analysis of LoRA’s individual rank.
@@
Table 7: L2 norm of $\Delta W$ signifies a larger update from the original weight $W$. STM shows a small model $\Delta W$, indicating STM training can occur with few changes on the model’s original capabilities.
Models tuned on MBPP Self-Output Rephrase Ground Truth STM + Ground Truth
Llama 3 8B Instruct 6.53 7.31 17.75 0.55
Gemma 2 IT 2B 4.03 5.78 5.69 0.45
High perplexity training leads to more degradation with similar weights updated
Although we found that directly training with high perplexity tokens leads to higher loss and more weights updated. Applying regularization techniques can reduce model weight changes to mitigate non-target task degradation. Hence, we apply alternative regularization techniques to fine-tune with high-perplexity ground truth data to mitigate non-target task performance degradation. Our comparative analysis encompasses three widely-adopted regularization approaches: Dropout [30], Weight Decay [17], and Kullback Leibler regularization [34]. These methods have demonstrated effectiveness in preventing models from overfitting. Table 8 shows that under similar model weight changes, models applied with STM still outperform the high perplexity data trained models with regularization, indicating the performance degradation results from high perplexity training instead of purely model overfitting. For the results of all regularization combinations we have experimented with, please refer to Appendix G.
Table 8: Comparison of STM and regularization strategies on the MBPP Ground Truth data using Llama3 8B Instruct. Hyperparameters are selected based on the performance or a similar L2 norm of $\Delta W$ to STM’s setup. STM consistently outperforms all regularization methods, suggesting that degradation is driven more by high-PPL training than by overfitting.
Regularization & Hyperparameter L2 norm of $\Delta W$ BWT (%) TI (%)
Weight decay 0 + dropout 0.05 0.7539 -9.24 -9.82
Weight decay 0.2 + dropout 0.3 0.7109 -3.50 -8.03
Weight decay 0.5 + dropout 0.3 0.5351 -11.15 2.86
KL $coef=1e^{-5}$ 2.3906 -7.96 -9.82
STM 0.5500 1.90 3.12
@@ 7 Related work
Instruction following [25, 5] is widely used to align responses of LLMs with target task values. Practically, it requires instruction tasks dataset to fine-tune LLMs, where each dataset consists of instructions and desired responses. Hence, the performance of instruction-tuning heavily relies on the quality of instruction data such as context richness [39, 35, 43]. In addition, Ghosh et al. [9] investigated the limitation of instruction-tuning of LLMs about the catastrophic forgetting from pattern-copying behaviors and hallucinations with Lora/full fine-tuning. These works give us a good start on how performance degrades in terms of models’ response behaviors and benchmarks when fine-tuning with instruction following dataset.
@@
As Wang et al. [35], Chung et al. [5] showed, LLMs break down easily after training with different tasks. Several remedies to the performance improvement focus on training data augmentation [36, 40]. For instance, [28] uses much larger LLMs (e.g., GPT-4 and Claude) to generate responses of questions as training labels, which improves the performance on both the target task and other non-target tasks. However, such a distillation method neglects the correctness of generated labels, so that more incorrect responses could be trained as the amount of generated data increases, and thus, using equal or smaller-sized models for distillation could be challenging. [41] prompts LLMs to simply rephrase the response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing the ground truth answer limits the output distribution and results in lower performance in our study. Furthermore, [11] exploits a base LLM as a judge to pick out answerable and unanswerable questions to compose a new training dataset to improve target task performance only on the QA/conversation dataset. Although using Mistral 7B Instruct to generate acceptable responses for answerable questions improves target and non-target tasks, it is rarely discussed that if the proposed method is applicable to different model sizes and series. Besides, using LLM as a judge could be a noise to the correctness of data, which is shown to be important for target task training in our study in Appendix F. Lastly, [19, 21] propose a Selective Language Modeling to score favorable tokens from training data, and pre-training on the tokens brings higher performance. However, the requirements for pre-training a reference model or preparing a high-quality dataset for a scoring model could be exhausting depending on task difficulty. In addition to STM’s applicability in Fine-Tuning, STM does not require pre-training for the selection of tokens at all, which brings efficiency and performance improvement at the same time.
@@ 8 Conclusion
In this work, we present a comprehensive empirical study of how fine-tuning with LLM-generated data enhances cross-domain generalization in instruction-following models. Our findings reveal that such data significantly reduces degradation on non-target tasks compared to traditional ground truth fine-tuning. This robustness is closely linked to the reduced presence of high perplexity tokens in input sequences, which we identify as a critical factor in preserving general capabilities. Building on this insight, we introduce the Selective Token Masking (STM) method—an efficient and model-agnostic strategy for achieving comparable improvements without relying on external model generations. Our extensive evaluations across diverse models, training strategies, and learning settings underscore the generality and effectiveness of low-perplexity training. These findings offer a loss- and perplexity-based perspective on fine-tuning robustness and point to a practical direction for future work: designing adaptive strategies that prioritize low-perplexity tokens or filter out harmful high-perplexity tokens. Such approaches may generalize across models and tasks, enabling efficient domain adaptation while preserving general capabilities.
9 Limitations
The effectiveness of token-wise masking on training models could relieve catastrophic forgetting and improve target task performance simply. Still, there are limitations not addressed: (1) Due to limitations of resources, we did not apply STM methods on LLMs larger than 10B. (2) Many more non-target tasks we have not evaluated yet, such as safety and instruction following benchmark. (3) Self-output data generation is costly. A trade-off between computing resources and performance can be further studied. (4) We did not consider how perplexity changes when task difficulty scales up. It may affect the robustness when training on harder tasks. Due to space and resource limitations, we leave the challenging issues as future work.
Appendices
Appendix A Main models chosen for comparison
We evaluate our method across different model scales and architectures using two publicly available instruction-tuned language models; here they are referred to as baseline models:
  • Gemma 2 IT 2B: A 2.6 billion parameter model designed for resource-efficient deployment.
  • Llama 3 8B Instruct: An 8 billion parameter model which has the best overall performance in all models.
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama 3 8B Insturct represent mid-scale models with different architectural innovations, while Gemma 2 IT 2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies. To further test the generalization of STM effectiveness on a brand new model we also choose three additional new models to test on target task MBPP and MATH, please refer to Appendix B
Appendix B Other LLM series performance
Similar to Table 2, we also have performances of additional 3 models as follows:
  • Mistral 7B-Instruct: A 7 billion parameter model featuring grouped-query attention and sliding window attention mechanisms.
  • Gemma 2 IT 9B Instruct: A 9 billion parameter model featuring as a larger sized LLM.
  • OLMo 2 7B Instruct: A 7 billion parameter model featuring as a much newer released model (2024 Dec.) that is different from currently experimented models.
We add the raw performance results of Mistral 7B-Instruct model of Self-Output, Rephrase, Ground Truth, and STM performances along with Gemma 2 IT 2B and Llama 3 8B Instruct as Table 10 shows: our conclusion that Self-Output and STM strategies for training data generation still holds for different model sizes and series. In addition, to demonstrates the generalization of our threshold choice settings, we also added the results of Gemma 2 IT 9B and OLMo 2 7B Instruct on MBPP target task with STM applied as Table 9 shows: STM’s optimal threshold selection on Gemma 2 IT 9B and OLMo 2 7B Instruct could reduce both non-target degradation and target-task improvement.
Table 9: STM performance of the new model on the MBPP dataset, compared to baseline fine-tuning.